OpenStudy (abbycross167):

Can someone please show me how to do this? I really need help... I will give a medal Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. y = -3x + 2

OpenStudy (abbycross167):

@Directrix Can you try to help me? Please?

OpenStudy (baru):

hi

OpenStudy (baru):

you will need graph paper

OpenStudy (abbycross167):

okay @baru I have a graph that I have to put it on

OpenStudy (baru):

first we will find "y intercept" y intercept is where the graph of the line cuts the y axis

OpenStudy (baru):

y=-3x + 2 to find y intercept; just substitute x=0 in the above equation what do you get?

OpenStudy (abbycross167):

-1?

OpenStudy (baru):

how did you get -1 i said substitute x=0 in the equation of the line, can you show your steps please

OpenStudy (abbycross167):

umm like this ~ -3+2

OpenStudy (baru):

y=-3x +2 substitute x=0 means:- write '0' in place of x y=-3(0) +2

OpenStudy (baru):

now can you find y?

OpenStudy (baru):

y=?

OpenStudy (abbycross167):

2? I'm sorry if I get it wrong.. I'm not very good at math

OpenStudy (baru):

thats ok :). it gets better with practice

OpenStudy (baru):

yes it is 2. but can you tell me how ?

OpenStudy (baru):

y= -3x + 2 here 3x actually means \(3 \times x\)

OpenStudy (abbycross167):

Umm I just did this~ -3(0)+2

OpenStudy (baru):

^you understood what i just said?

OpenStudy (abbycross167):

yes sir/ma'am

OpenStudy (baru):

lol... i'm a guy, but dont cal me sir

OpenStudy (baru):

when i say substitute x=0, you just have to write 0 in place of x y= -3x +2 y=-3(0) +2 y=0 +2 (because 0 multiplied with any number is zero) y=2

OpenStudy (baru):

followed?

OpenStudy (abbycross167):

yes

OpenStudy (baru):

ok, now y=2 is the y intercept in other words, its the point where the line cuts the y axis

OpenStudy (baru):

on your graph sheet, mark the x axis and y axis

OpenStudy (baru):

@abbycross167

OpenStudy (abbycross167):

Ok I'm sorry I have slow WiFi. y= 2 and x=0?

OpenStudy (baru):

yes exactly, mark that point in yellow

OpenStudy (baru):

done with that?

OpenStudy (abbycross167):

yes sir

OpenStudy (baru):

XD no 'sir'

OpenStudy (abbycross167):

haha sorry

OpenStudy (baru):



OpenStudy (baru):



OpenStudy (baru):

i've marked the point with a circle, does your graph look similar

OpenStudy (baru):

?

OpenStudy (abbycross167):

yes

OpenStudy (baru):

ok, now we will find 'x intercept' y=-3x + 2 so substitute y=0 what do you get?

OpenStudy (abbycross167):

is 'x' still 0?

OpenStudy (baru):

no, to find x intercept, we start fresh, so leave 'x' as x

OpenStudy (baru):

lets go step by step substitute y=0 in the equation, write what you get

OpenStudy (abbycross167):

0=-3x+2?

OpenStudy (baru):

yes perfect now, we need to solve this equation and find 'x'

OpenStudy (baru):

do you know how to do that?

OpenStudy (abbycross167):

umm not really

OpenStudy (baru):

ok, the rule for solving equations is, you can do anything you want; add numbers or subtract or multiply or divide, but what ever you do, you have to do it to BOTH sides of the equation

OpenStudy (baru):

0=-3x + 2 first, subtract 2 from BOTH SIDES of the equation write what you get

OpenStudy (abbycross167):

-3-2+2-2=-5?

OpenStudy (baru):

no, leave everything as is, do not remove x 0 = -3x + 2 subtract 2 from both sides 0\(-2\)=-3x +2 \(-2\)

OpenStudy (baru):

do you understand what i just did?

OpenStudy (abbycross167):

umm I think so... Is it -3?

OpenStudy (baru):

no... we are not there yet 0-2= -3x + 2 -2 can you simplify this a little there is 2-2 on the right side of the equation

OpenStudy (abbycross167):

2-2=0... so that just leaves me with -3x?

OpenStudy (baru):

correct!!, you are left with just -3x ON THE RIGHT SIDE OF THE EQUATION what is on the left side?

OpenStudy (abbycross167):

2? or -2?

OpenStudy (baru):

0-2=-2 so now can you write the whole equation after simplifying 0-2 and 2-2 ?

OpenStudy (abbycross167):

I'll try

OpenStudy (baru):

like you said ON THE RIGHT SIDE we are left with just -3x ON THE LEFT SIDE we have -2 so what is the equation now?

OpenStudy (abbycross167):

would it just be -3x-2 ?

OpenStudy (baru):

0-2= -3x + 2 -2 i am just asking you to simplify this 0-2 is -2 2-2 is 0 -2=-3x +0 -2= -3x

OpenStudy (abbycross167):

oo ok

OpenStudy (baru):

so now we have -2 = -3x divide BOTH SIDES OF THE EQUATION BY -3 what do you get?

OpenStudy (abbycross167):

-2 divided by -3=1.5 -3 divided by -3=1 ?

OpenStudy (baru):

yes!!!

OpenStudy (baru):

very good :) now write that in the the equation

OpenStudy (abbycross167):

1.5=1x?

OpenStudy (baru):

yes, excellent! you are catching on

OpenStudy (baru):

you know that 1 multiplied with any number is the same number so 1x is just x

OpenStudy (baru):

oh sorry, i didnt notice, -2 divided by -3 is not 1.5, its 0.67 its just a minor mistake...but your step are correct

OpenStudy (baru):

so can you correct that, and use what i just said about 1x and re-write the equation

OpenStudy (abbycross167):

ok so the equation would be 0.67=1x?

OpenStudy (baru):

yes and 1x=x so the equation is 0.67=x congratulations, you have solved the equation and found x. :)